# FOCUS on Contamination: A Geospatial Deep Learning Framework with a Noise-Aware Loss for Surface Water PFAS Prediction

### v3

@@ Results
Accuracy 80 $\pm$ 2 85 $\pm$ 1 92 $\pm$ 1 96 $\pm$ 1 90 $\pm$ 2 91 $\pm$ 1
IoU 64 $\pm$ 2 67 $\pm$ 1 65 $\pm$ 2 73 $\pm$ 1 75 $\pm$ 2 77 $\pm$ 1
F-score 77 $\pm$ 1 79 $\pm$ 1 76 $\pm$ 2 82 $\pm$ 1 85 $\pm$ 1 87 $\pm$ 1
Precision 78 $\pm$ 2 87 $\pm$ 1 70 $\pm$ 3 83 $\pm$ 2 81 $\pm$ 2 83 $\pm$ 1
Recall 76 $\pm$ 2 77 $\pm$ 1 96 $\pm$ 1 91 $\pm$ 2 94 $\pm$ 1 92 $\pm$ 1
Table 5: Cross-state generalization results ($\text{mean}\pm\text{std}$). Bold values denote the best result per year.
To further assess the generalizability of our model to geographically unseen regions, we performed a spatial cross-validation experiment by partitioning data based on U.S. states. Specifically, we selected three distinct sets of five states each to serve as held-out test sets. The remaining states in each case were used exclusively for training. This ensures that no spatial overlap exists between training and testing regions, mimicking the real-world scenario where models are deployed in previously unsampled areas. FOCUS again consistently outperforms focal loss alone across all years, highlighting its ability to generalize spatially to unseen regions and its robustness in distinguishing contaminated areas even under strong distribution shifts.
To assess the spatial robustness of our model, we also performed a consistency analysis across all 49 continental U.S. states. This involved comparing predictions over overlapping regions from multiple randomly sampled patches per state, with varying degrees of overlap (56×56 and 156×156 pixels). Agreement was measured as the proportion of matching predicted labels across overlaps. We found high agreement scores (above 93% on average), indicating that our model is locally consistent in its predictions. Full details are provided in the supplement.
Results on model calibration (i.e., Expected Calibration Error (ECE) (Guo et al. 2017)) are provided in the supplement, showing low ECE values ($\leq 0.1$) and indicating that predicted probabilities are well-aligned with observed outcomes.
Collaboration with Stakeholders
To build this model, and ultimately deploy it in a way that meets real-world needs, we are collaborating with a nonprofit environmental organization. This partnership has helped shape both the technical and practical framing of our work. We meet biweekly to discuss results, align on next steps, and ensure that public-facing outputs, such as PFAS contamination maps, are communicated transparently and responsibly. Recognizing the ethical implications of sharing such information, including for public health, environmental justice, and even property values, we are also currently engaging with an advisory group with representatives from other non-profits and agencies. This process will help tailor outputs to community needs and support responsible use in remediation and policy decisions.
Conclusion and Future Work
PFAS contamination is a global environmental issue due to its persistence in water, soil, and wildlife. While lab-based methods like LC-MS offer high accuracy, they are expensive and slow. Alternative approaches, such as spatial interpolation, remote sensing, ML methods, and hydrological models, provide insights, but struggle with feature engineering and loss of spatial context.
Our geospatial DL framework, FOCUS, directly processes multi-channel raster data and integrates a novel noise-aware loss, improving detection performance and preserving spatial dependencies, making it a promising tool for large-scale PFAS contamination mapping that can help identify high-risk areas.
@@
Future work will explore uncertainty estimation to guide targeted sampling and enhance prediction robustness. We also aim to incorporate temporal modeling using multi-year data to track and forecast evolving PFAS patterns. Finally, we plan to build on public contamination maps, such as those from the Environmental Working Group (Environmental Working Group 2024) and (U.S. Geological Survey 2023), by deploying our AI-based framework responsibly.
Supplement
To help navigate the supplementary material, Table 6 provides an overview linking each supplementary subsection to the corresponding sections in the main paper where they are mentioned or expanded upon.
Supplement Subsection Referenced in Main Paper Section(s)
@@
B. Noise Mask Weights Ablation Section 3.1
@@ H. FOCUS Per-Class Results -
I. Ternary Classification of Hazard Levels Section 3.1
@@
K. Distance-to-Industry Rasters. Section 3.1
@@
Table 6: Mapping between supplementary sections and corresponding references in the main paper.
A Model Calibration
We evaluated model calibration using ECE as introduced in Section 4.4, which quantifies the discrepancy between predicted probabilities and observed outcomes where lower scores indicate better calibration. Table 7 shows the ECE scores for our model trained with FOCUS loss for the years 2008, 2019, and 2022.
@@ A Model Calibration
Table 7: ECE Scores for Models Trained with FOCUS Loss Across Different Years.
@@
B Noise Mask Weights Ablation
Our ablation study on the 2008 PFAS dataset evaluated a comprehensive set of 24 noise mask weight configurations through a grid search over four input types: dischargers, land cover, flow direction, and distance to sample points as mentioned in Section 3.1. For each configuration, we varied the percentage weights assigned to these four sources of uncertainty and measured segmentation performance across multiple metrics. The best-performing configuration assigns 40% weight to dischargers, 20% to land cover, 10% to sample distance, and 30% to flow direction. This setup achieved the highest scores in accuracy, IoU, F-score, and precision, even though a few alternate combinations yielded marginally higher recall. These results align with domain expert guidance that emphasizes the importance of industrial activity in driving contamination patterns. Full performance results for all configurations are reported in Table 8, demonstrating that our selected weighting scheme most effectively balances the different environmental signals for robust PFAS contamination prediction.
@@ B Noise Mask Weights Ablation
40 20 30 10 65 77 83 73 88
40 30 20 10 44 52 67 50 76
40 10 30 20 48 60 68 57 80
40 10 20 30 43 51 52 51 78
40 30 10 20 49 59 63 58 82
40 20 10 30 47 55 67 54 84
30 40 10 20 56 69 93 61 80
30 40 20 10 44 58 58 64 69
30 10 40 20 46 56 59 55 76
30 10 20 40 46 60 59 65 71
30 20 40 40 54 67 67 67 82
30 20 10 40 51 63 63 65 78
30 10 30 30 42 46 42 50 84
10 40 30 20 59 71 73 69 85
10 40 20 40 47 59 58 59 76
10 30 40 20 43 51 52 51 78
10 30 20 40 41 57 62 73 62
10 20 40 10 49 63 63 70 73
10 20 30 40 44 55 55 55 76
20 40 30 40 51 67 68 77 69
20 40 10 30 51 61 68 59 84
20 30 40 10 47 62 63 73 69
20 30 20 30 45 60 64 75 65
20 10 40 30 35 52 64 64 47
Table 8: Performance across all noise mask weight configurations using the 2008 PFAS dataset.
C Kriging
We employed Kriging as a spatial interpolation technique to estimate PFAS, as mentioned in Section 3.3. First, we examined the spatial structure of our data by calculating an empirical semivariogram. Using latitude and longitude coordinates alongside the binary target variable, we computed pairwise distances and differences to produce a semivariogram plot, enabling us to observe how variance changed with increasing distance. This step informed the selection of appropriate variogram models (e.g., linear or spherical) for Kriging. Based on this analysis, we selected a spherical variogram model, which best captured the spatial dependencies observed in our data.
@@ C Kriging
D Pollutant Transport Simulation
As outlined in Section 3.3, instead of running a simulation like SWAT directly, which is designed for general watershed hydrology modeling but lacks established parameterization for PFAS transport, we implemented a streamlined Python-based workflow more tailored to PFAS-related surface water contamination. This approach allowed us to integrate domain-relevant features without relying on assumptions unsupported by current PFAS science.
Figure 8: Example binary output from the pollutant transport simulation: 0 represents low contamination, and 1 represents high contamination. The simulation approximates the distribution of PFAS contamination based on hydrological and environmental parameters, providing a practical alternative to full-scale SWAT simulations.
@@ D Pollutant Transport Simulation
E Consistency Analysis
To evaluate local spatial consistency as highlighted in Section 4.4, we applied a robustness test across all 49 continental U.S. states. For each state, we extracted two overlapping image patches; one with a 56×56 pixel overlap and another with 156×156, using three different random locations to ensure stability. We then compared the model’s predictions over the overlapping regions by computing pixel-wise agreement. Agreement is the proportion of matching predicted labels across overlapping pixels between two patches:
$\text{Agreement}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(y_{i}^{(1)}=y_{i}^{(2)}),$ (5)
@@ E Consistency Analysis
Table 9: Spatial consistency measured as prediction agreement over overlapping patch regions.
F Evaluating the Impact of Noise Masks on Segmentation of Land Cover Types
@@
Results
Below are the macro-averaged results showing the performance metrics for the Evergreen Forests vs Barren Land and Deciduous Forests vs Water pairs, both with and without noise masks. These LC image patches are aligned with the same spatial footprint as the PFAS patches to ensure consistency in coverage. The evaluation is conducted on at most 2 pixels per image patch, simulating sparse ground truth availability similar to the PFAS setting. We see that including noise masks is useful in this setting as well.
@@ Results
Table 10: Performance Comparison for Evergreen Forests vs Barren Land: Simulated Uncertainty in PFAS Pseudolabeling
@@
Table 11: Performance Comparison for Deciduous Forests vs Water: Simulated Uncertainty in PFAS Pseudolabeling
G Dataset
As introduced in Section 3.3 and expanded upon here, the dataset was curated to ensure that patches included in the training and testing sets were geographically disjoint, avoiding any overlap.
@@ G Dataset
Figure 9: Example of overlapping patch assignment. The figure illustrates two slightly overlapping patches overlaid on the U.S. map, each containing distinct sample points. To maintain disjoint training and testing sets, both patches are assigned to the same set (either training or testing), ensuring no part of a test patch overlaps with any patch seen during training.
Additionally, Table 12 expands on the table in Section 3.1 to include class distribution. There is a significant class imbalance that we observe in the real world: unfortunately, a majority of water bodies have PFAS contamination that exceed health guidelines, as found in prior scientific work on PFAS prediction as well  [^2] See: PFAS pollute 83% of U.S. waterways, E&E News by Politico (2023). Available at: https://www.eenews.net/articles/pfas-pollute-83-of-u-s-waterways/ . To affirm that there is in fact signal to our predictions, we provide class-specific results in Section H. For instance, we find approximately 83% F-score for class 0 for the year 2024, showing we are truly predicting class 0, and better than random. We also add a ternary classification in Section I to provide additional levels of risk.
@@
Fish Train/Test EPA (U.S. EPA, OW 2015; US EPA 2024) 866 89.5% / 10.5%
Fish Validation MPART (Michigan Department of Environment, Great Lakes, and Energy 2025) 114 77.2% / 22.8%
Water Train/Test MPART (Michigan Department of Environment, Great Lakes, and Energy 2023) 293 66.9% / 33.1%
@@
Table 12: Datasets used for training, testing, and real-world validation, including total size and class balance.
H FOCUS Per-Class Results
@@
Accuracy 88 $\pm$ 1 88 $\pm$ 1 89 $\pm$ 1 89 $\pm$ 1 86 $\pm$ 2 86 $\pm$ 2 86 $\pm$ 2 86 $\pm$ 1
IoU 43 $\pm$ 2 87 $\pm$ 2 53 $\pm$ 2 88 $\pm$ 2 52 $\pm$ 3 83 $\pm$ 2 71 $\pm$ 2 78 $\pm$ 2
F-score 60 $\pm$ 2 93 $\pm$ 2 70 $\pm$ 2 93 $\pm$ 2 68 $\pm$ 3 91 $\pm$ 2 83 $\pm$ 2 88 $\pm$ 2
Precision 75 $\pm$ 3 90 $\pm$ 2 57 $\pm$ 2 98 $\pm$ 2 52 $\pm$ 3 99 $\pm$ 1 99 $\pm$ 1 78 $\pm$ 2
Recall 50 $\pm$ 2 96 $\pm$ 2 89 $\pm$ 1 89 $\pm$ 1 99 $\pm$ 1 83 $\pm$ 2 71 $\pm$ 1 99 $\pm$ 1
Table 13: Per-class performance of our model using FOCUS across four years (MPART surface water results at 2024). Results averaged over 3 random seeds.
I Ternary Classification of Hazard Levels
To enable finer-grained prioritization of contamination severity, we extended our binary classification framework to a ternary setting, as highlighted in Section 3.1. Specifically, each fish tissue sample was assigned to one of three classes based on its hazard index (HI), defined as the ratio of measured PFAS concentration to the EPA’s reference dose:

### v4

@@ 5.4. FOCUS has Robust Results
Accuracy 58 $\pm$ 2 70 $\pm$ 1 76 $\pm$ 1 89 $\pm$ 1 78 $\pm$ 2 92 $\pm$ 1
IoU 34 $\pm$ 2 48 $\pm$ 1 40 $\pm$ 2 55 $\pm$ 1 56 $\pm$ 2 75 $\pm$ 1
F-score 47 $\pm$ 1 63 $\pm$ 1 46 $\pm$ 2 61 $\pm$ 1 68 $\pm$ 1 85 $\pm$ 1
Precision 49 $\pm$ 2 62 $\pm$ 1 49 $\pm$ 3 59 $\pm$ 2 69 $\pm$ 2 82 $\pm$ 1
Recall 49 $\pm$ 2 72 $\pm$ 1 50 $\pm$ 1 71 $\pm$ 2 85 $\pm$ 1 91 $\pm$ 1
6. Scientific Impact and Interdisciplinary Considerations
This work advances PFAS science by linking large-scale environmental context to surface-water contamination risk and its implications for bioaccumulation. By integrating industrial dischargers, land cover, and hydrological transport with sparse measurements, our framework produces spatially consistent risk maps that help identify regions likely to be under-sampled, supporting targeted follow-up testing and more systematic risk assessment. At a broader scale, such modeling enables more consistent, nationwide perspectives on contamination risk, complementing existing, state-level fish advisory programs.
Beyond static risk mapping, our interface supports scientific exploration and hypothesis generation. From a scientific perspective, while FOCUS does not yet perform source attribution, the interface enables exploratory analysis by overlaying predicted risk with known discharger locations and hydrological context. Regions exhibiting high predicted risk despite few nearby known sources, such as the highlighted example in Fig 9, have also been noted by domain experts as areas of unexplained contamination, suggesting potential unknown, legacy, or poorly characterized sources and motivating targeted follow-up investigation.
The need for AI/ML in this domain is driven by severe data sparsity and fragmentation. PFAS comprise thousands of compounds with diverse sources and uses, yet environmental measurements—especially in surface waters and fish—remain extremely limited in space and time. Traditional monitoring and mechanistic modeling alone cannot provide comprehensive coverage, motivating learning-based approaches that integrate heterogeneous geospatial data and extrapolate risk patterns under uncertainty.
Applying AI in this setting also raises scientific and societal challenges, including interpretability, uncertainty communication, and regulatory alignment. Our approach addresses these concerns by explicitly modeling spatially structured label noise and incorporating physically informed priors, providing more calibrated and transparent predictions than standard black-box models.
7. FOCUS Web Map Interface
To support real-world use and stakeholder engagement, we have developed a web-based, interactive PFAS risk mapping interface with policymakers and domain stakeholders that visualizes model predictions at national scale (Fig. 9). The interface displays predicted PFAS contamination risk across surface waters and fish tissue using discrete risk categories (e.g., low, medium, high), enabling intuitive exploration by non-technical users.
Figure 9. Web interface for exploring FOCUS predictions. The zoomed region shows high predicted risk despite few nearby known dischargers (shown as point overlays).
The map integrates FOCUS predictions with contextual geospatial layers, including surface water extents, land cover, hydrological features, and known discharge locations. Users can toggle layers, zoom to regions of interest, and inspect spatial patterns across years and data modalities. The interface was developed in collaboration with policymakers and domain stakeholders to ensure interpretability, transparency, and responsible communication of model outputs.
This deployment illustrates how noise-aware geospatial predictions can be translated into actionable decision-support tools, while acknowledging uncertainty and the limitations of sparse environmental sampling.
8. Conclusion and Future Work
We introduced FOCUS, a geospatial deep learning framework that integrates multi-channel raster inputs with a principled noise-aware learning objective designed for spatially structured label noise arising from point-to-pixel supervision. By combining hydrology-informed correctness priors with a noise-robust loss, FOCUS improves predictive performance while preserving spatial coherence, enabling large-scale PFAS contamination screening.
@@ 8. Conclusion and Future Work
Future work will explore predictive uncertainty quantification to guide targeted sampling and enhance prediction robustness. We also aim to incorporate temporal modeling using multi-year data to track and forecast evolving PFAS patterns, as longitudinal measurements become denser and domain understanding enables more physics-informed supervision and priors. Finally, we plan to build on public contamination maps (e.g., EWG (Environmental Working Group, 2024), (U.S. Geological Survey, 2023)) by deploying our framework responsibly.
9. Limitations and Ethical Considerations
As with any large-scale environmental modeling effort, FOCUS performance is ultimately bounded by the availability and quality of underlying data. Additionally, we consider a static, one-shot prediction setting; coupling FOCUS with active or adaptive sampling is left to future work. Furthermore, the current model does not yet attribute predicted risk to specific PFAS compounds and sources, though we are actively developing extensions in this direction. Accordingly, predictions should be interpreted as screening-level risk signals to prioritize follow-up sampling rather than definitive contamination assessments.
Given the ethical implications of releasing public-facing PFAS risk maps—including impacts on public health, environmental justice, and property values—we engage an advisory group with representatives from nonprofits and agencies (e.g., HRWC). This collaboration helps tailor outputs to community needs, communicate uncertainty, and support responsible use in remediation and policy decisions.
appendix Subsection Referenced in Main Paper Section(s)
@@ 9. Limitations and Ethical Considerations
B. Physically-Informed Pixel Confidence $M_{i}$ Details Section 3.4
@@ H. FOCUS Per-Class Results -
I. Ternary Classification of Hazard Levels Section 3.1.1
@@
K. Distance-to-Industry Rasters. Section 3.1.2
@@
O. FOCUS Loss Derivation and Assumptions Section 3.5
P. Region-Level Risk Signals and Bioaccumulation Evidence -
Appendix A Model Calibration
We evaluated model calibration using ECE, which quantifies the discrepancy between predicted probabilities and observed outcomes where lower scores indicate better calibration. Table 7 shows the ECE scores for our model trained with FOCUS loss for the years 2008, 2019, and 2022.
Table 7. ECE Scores for Models Trained with FOCUS Loss Across Different Years.
@@ Appendix A Model Calibration
Appendix B Physically-Informed Pixel Confidence $M_{i}$ Details
we generate a domain expert-informed probability of the label correctness for each surface water pixel $i$:
(5) $\,+\alpha_{4}\cdot p_{\text{downstream}}$
  • Proximity to PFAS Dischargers ($p_{\text{dischargers}}$) Proximity to known PFAS dischargers is a strong indicator of contamination risk, as shown in scientific analyses  [^3] PFAS dischargers. If a pixel’s label is 1, we apply an exponential decay function of distance, assigning higher probabilities to pixels closer to known dischargers. For pixels labeled 0, we invert this logic so that being near a discharger lowers confidence in the 0 label.
  • Land Cover Type ($p_{\text{landcover}}$) For pixels labeled 1, surrounding urban or built-up areas received higher probabilities, consistent with research linking PFAS to industrial zones   [^4] Landcover. Conversely, for 0-labeled pixels, surrounding undeveloped or forested land cover increased confidence in low contamination.
  • Proximity to Other Sample Points ($p_{\text{sample\_dist}}$) Pixels closer to ground truth sample points were assigned higher probabilities using an exponential decay function based on distance to each sample point in the raster.
  • Downstream Flow ($p_{\text{downstream}}$) PFAS may also travel downstream in aquatic systems, making hydrologic flow another critical factor in contamination spread  [^5] Downstream Flow. If a pixel lies downstream of a sample point labeled 1, its probability for label 1 increases, and similarly for label 0. This flow direction channel was generated using ArcGIS Pro  [^6] ArcGIS Pro based on digital elevation models (DEMs)  [^7] DEM.
Full performance results for all configurations are reported in Table 8, demonstrating that our selected weighting scheme most effectively balances the different environmental signals for robust PFAS contamination prediction.
Table 8. Performance across all noise mask weight configurations using the 2008 PFAS dataset.
@@ Appendix B Physically-Informed Pixel Confidence $M_{i}$ Details
40 20 30 10 48 61 61 64 73
40 30 20 10 44 55 58 52 70
40 10 30 20 46 59 60 56 72
40 10 20 30 42 53 55 51 69
40 30 10 20 45 57 59 54 71
40 20 10 30 43 54 57 52 70
30 40 10 20 46 58 60 55 72
30 40 20 10 43 56 57 53 68
30 10 40 20 44 55 56 53 69
30 10 20 40 45 57 58 55 70
30 20 40 40 47 62 60 58 72
30 20 10 40 45 58 59 56 71
30 10 30 30 41 52 54 49 69
10 40 30 20 47 60 60 58 72
10 40 20 40 44 56 57 54 69
10 30 40 20 42 53 55 51 70
10 30 20 40 43 55 56 52 68
10 20 40 10 46 59 60 57 72
10 20 30 40 44 56 57 53 70
20 40 30 40 45 58 59 56 69
20 40 10 30 46 59 60 57 72
20 30 40 10 44 57 58 55 69
20 30 20 30 43 56 57 54 68
20 10 40 30 40 52 54 50 66
Appendix C Kriging
We employed Kriging as a spatial interpolation technique to estimate PFAS. First, we examined the spatial structure of our data by calculating an empirical semivariogram. Using latitude and longitude coordinates alongside the binary target variable, we computed pairwise distances and differences to produce a semivariogram plot, enabling us to observe how variance changed with increasing distance. This step informed the selection of appropriate variogram models (e.g., linear or spherical) for Kriging. Based on this analysis, we selected a spherical variogram model, which best captured the spatial dependencies observed in our data.
@@ Appendix C Kriging
Appendix D Pollutant Transport Simulation
Instead of running a simulation like SWAT directly, which is designed for general watershed hydrology modeling but lacks established parameterization for PFAS transport, we implemented a streamlined Python-based workflow more tailored to PFAS-related surface water contamination. This approach allowed us to integrate domain-relevant features without relying on assumptions unsupported by current PFAS science.
Figure 10. Example binary output from the pollutant transport simulation: 0 represents low contamination, and 1 represents high contamination. The simulation approximates the distribution of PFAS contamination based on hydrological and environmental parameters, providing a practical alternative to full-scale SWAT simulations.
@@ Appendix D Pollutant Transport Simulation
Appendix E Consistency Analysis
To evaluate local spatial consistency, we applied a robustness test across all 49 continental U.S. states. For each state, we extracted two overlapping image patches; one with a 56×56 pixel overlap and another with 156×156, using three different random locations to ensure stability. We then compared the model’s predictions over the overlapping regions by computing pixel-wise agreement. Agreement is the proportion of matching predicted labels across overlapping pixels between two patches:
(6) $\text{Agreement}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(y_{i}^{(1)}=y_{i}^{(2)}),$
@@ Appendix E Consistency Analysis
Table 9. Spatial consistency measured as prediction agreement over overlapping patch regions.
@@
Appendix F Evaluating the Impact of Noise Masks on Segmentation of Land Cover Types
@@
Results Below are the macro-averaged results showing the performance metrics for the Evergreen Forests vs Barren Land and Deciduous Forests vs Water pairs, both with and without noise masks. These LC image patches are aligned with the same spatial footprint as the PFAS patches to ensure consistency in coverage. The evaluation is conducted on at most 2 pixels per image patch, simulating sparse ground truth availability similar to the PFAS setting. We see that including noise masks is useful in this setting as well.
Table 10. Performance Comparison for Evergreen Forests vs Barren Land: Simulated Uncertainty in PFAS Pseudolabeling
@@
Table 11. Performance Comparison for Deciduous Forests vs Water: Simulated Uncertainty in PFAS Pseudolabeling
@@
Appendix G Dataset
The dataset was curated to ensure that patches included in the training and testing sets were geographically disjoint, avoiding any overlap.
@@ Appendix G Dataset
Figure 11. Example of overlapping patch assignment. The figure illustrates two slightly overlapping patches overlaid on the U.S. map, each containing distinct sample points. To maintain disjoint training and testing sets, both patches are assigned to the same set (either training or testing), ensuring no part of a test patch overlaps with any patch seen during training.
Additionally, Table 12 expands on the table in Section 3.5.1 to include class distribution. There is a significant class imbalance that we observe in the real world: unfortunately, a majority of water bodies have PFAS contamination that exceed health guidelines, as found in prior scientific work on PFAS prediction as well  [^8] See: PFAS pollute 83% of U.S. waterways, E&E News by Politico (2023). Available at: https://www.eenews.net/articles/pfas-pollute-83-of-u-s-waterways/ . To affirm that there is in fact signal to our predictions, we provide class-specific results in Section H. For instance, we find approximately 80% F-score for class 0 for the year 2024, showing we are truly predicting class 0, and better than random. We also add a ternary classification in Section I to provide additional levels of risk.
Table 12. Datasets used for training, testing, and real-world validation, including total size and class balance.
@@
Fish Train/Test EPA (U.S. EPA, OW, 2015; US EPA, 2024) 866 89.5% / 10.5%
Fish Validation MPART (Michigan Department of Environment, Great Lakes, and Energy, 2025) 114 77.2% / 22.8%
Water Train/Test MPART (Michigan Department of Environment, Great Lakes, and Energy, 2023) 293 66.9% / 33.1%
@@
Appendix H FOCUS Per-Class Results
@@
Table 13. Per-class performance of our model using FOCUS across four years (MPART surface water results at 2024). Results averaged over 3 random seeds.
@@
Accuracy 73 $\pm$ 1 73 $\pm$ 1 82 $\pm$ 1 82 $\pm$ 1 88 $\pm$ 2 88 $\pm$ 2 79 $\pm$ 2 79 $\pm$ 1
IoU 25 $\pm$ 2 70 $\pm$ 2 40 $\pm$ 2 79 $\pm$ 2 47 $\pm$ 3 87 $\pm$ 2 67 $\pm$ 2 63 $\pm$ 2
F-score 40 $\pm$ 2 82 $\pm$ 2 57 $\pm$ 2 88 $\pm$ 2 64 $\pm$ 3 93 $\pm$ 2 80 $\pm$ 2 77 $\pm$ 2
Precision 33 $\pm$ 3 88 $\pm$ 2 42 $\pm$ 2 98 $\pm$ 2 60 $\pm$ 3 94 $\pm$ 1 75 $\pm$ 1 83 $\pm$ 2
Recall 50 $\pm$ 2 78 $\pm$ 2 89 $\pm$ 1 80 $\pm$ 1 69 $\pm$ 1 92 $\pm$ 2 86 $\pm$ 1 71 $\pm$ 1
Appendix I Ternary Classification of Hazard Levels
To enable finer-grained prioritization of contamination severity, we extended our binary classification framework to a ternary setting. Specifically, each fish tissue sample was assigned to one of three classes based on its hazard index (HI), defined as the ratio of measured PFAS concentration to the EPA’s reference dose: